given that tan y=√2, where x and y are acute angles. find cos y?


1
Expert's answer
2020-10-15T18:45:28-0400

tan y = "\sqrt{2}" Is also the same as,


tan y = "\frac{\sqrt{2}}{1}"

Hence

"\sqrt{2}" Is the opposite while "1" is the adjacent.



Hence

Hypotenuse(H) = "\sqrt{(\sqrt{2})^2+1^2}"

H = "\sqrt{3}"

Therefore cos y = "\frac{Adjacent}{Hypotenuse}"

Cos y = "\frac{1}{\sqrt{3}}"